Please answer the following question correctly please:

Problem 2-3 Compute the multifactor productivity measure for each of the weeks shown for production of chocolate bars. Assume 40-hour weeks and an hourly wage of $16. Overhead is 1.5 times weekly labor cost. Material cost is $9 per pound. (Round your answers to 2 decimal places.) Output (units) 32,000 37,000 28,000 38,000 Week Workers Material (lbs) 470 520 480 580 3 Week MFP (output /Total Cost)
Answer #1
Calculations A B C = B x 40 D = C x $16 E = D x 1.5 F G = F x $9 H = D+E+G I = A / H
Week Output Workers Hours worked Wages Overhead Material quantity Material cost Total cost MFP
1    32,000 5 200           3,200          4,800 470         4,230      12,230 2.62
2    37,000 6 240           3,840          5,760 520         4,680      14,280 2.59
3    28,000 6 240           3,840          5,760 480         4,320      13,920 2.01
4    38,000 7 280           4,480          6,720 580         5,220      16,420 2.31